Clabber Girl (CG) forecasts dividend growth of 10%, 20% and 14% for the next three years before settling to a long term growth rate of 5%. If the last dividend paid by CG was $5.2, what is the price for one share of CG. Additional information: B = 1.8, Market Risk Premium = 7%, Risk-free rate = 3%.
Answer #1

Required Return, rs = Risk-free Rate + Beta * Market Risk Premium
Required Return, rs = 3.00% + 1.80 * 7.00%
Required Return, rs = 15.60%

Last Dividend, D0 = $5.20

Growth rate for next 3 years are 10%, 20% and 14% and a constant growth rate (g) of 5%

D1 = $5.20 * 1.10 = $5.72
D2 = $5.72 * 1.20 = $6.864
D3 = $6.864 * 1.14 = $7.82496
D4 = $7.82496 * 1.05 = $8.216208

P3 = D4 / (rs - g)
P3 = $8.216208 / (0.1560 - 0.05)
P3 = $8.216208 / 0.1060
P3 = $77.5114

P0 = $5.72/1.1560 + $6.864/1.1560^2 + $7.82496/1.1560^3 + $77.5114/1.1560^3
P0 = $65.33

So, current stock price is $65.33